Prev Next

SW&IT Meeting Minutes 2018-07-30

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • Eng. manager on-site interview
  • new Telescope Eng - Peter Gray
  • new OSA - Josh Williams
  • PP2019 draft
  • SW Eng. recruitment
  • UCAP
    • State wide, AZ has adopted
    • To benchmark jobs, align salaries, to ID career paths
    • Titles will be per \x93market-based job functions and families
    • Salary and promotions will be defined by the job function/family
    • Employment will be \x93at will\x94

Previous Action Items

  • None

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • revision control workflow (discussion) [MB]
  • coding standards, indent tools (discussion) [MB]
  • ops-support oriented `single` wiki page (needs contributors) [MB]
  • circulate SSD plan (Leroy and Chris in the loop) [MB]
  • SH tech talk #3 [MB]
  • international travel 30days notice UA policy [MB]
  • Veeam HostCPUUsageAlarm update from David [MB]
  • git migration, next? [MB]

This week in github

Team reports

AC

  • Jenkins TAN RPM build is failing
  • What's the best way to exchange data between TCS and ARGOS swing arm script - some type of parser or json?

DE

Last week:
  • CpuUsageAlarm was a transient; it was cleared by the time I checked VMWare
  • Found fix for AGW computer requiring human intervention after unceremonious shutdown; now we won't need physical access or a KVM to it if the power is pulled
  • Started research into VMWare upgrades
  • Made historical backup of Doug's laptop hard drive
  • Reimaged that laptop into a SWG general use Linux machine
  • Started modifying IT email template to remove obviously redundant information (LBTO, etc). Should we ask the users what is actually useless before continuing?
  • Wrangling Dell on Transferring machines from Germany to us in order to refresh service contracts. Heard back to day that the transfer is actually complete this time.
This week:
  • Sort out VMWare upgrade whys and wherefores
  • Change Instruments Issue Type to IT internal lists? Do we consider Robert the ultimate authority for the list membership? Should Robert be sole Instrument Manager?

IS

This week:
  • Last Monday:
    • field AF LUCI test/dev python script. I still have some improvements to do that I couldn't finish on monday, namely using python canvas instead of the matplotlib GUIs.
  • Vacation for rest of the week

This week:
  • will be LUCI work:
    • .- focus offsets in a per grating and camera adjustments
    • .- AFC work

KS

Last Week - one half day
  • OAC HK checkout of a couple issues. Meeting with Dan and Matthieu about alarm conditions, client program conflicts with HK loop and how it should be fixed.

This Week
  • OVMS kernel update status - test after they get it finally booting?
  • OAC HK modifications
    • initial state of AGWs should be ignored for alarming purposes
    • HK disconnected should not alarm unless it happens for some time - disconnect and reconnect should be ok
    • "OK" message on return from 8051 sometimes has an extra first character, so OAC complains the command didn't work
    • mutex the use of the HK port by loop and clients instead of letting it fail in either case
    • create ALH config file for new OAC alarms
  • Add new GCS DD entries for all the 8051 HK data for 2018B release

MB

  • management OC, (no WPC)
  • eng. manager interview
  • catching up 500+ email pending inbox (Alfio jumbo network, Florian LN/docker, Rick MODS github, Marco Shark-nir interface, etc.)
  • review 30 applicants, shortlist 3, 10 under consideration, follow up aas job register
  • OVMS+TT follow up with martin/jorge-uwe @ ISYS, review go-nogo when Al comes back
  • follow up OAC/GCS work with Kellee + Dan
  • follow up OVMS uei upgrade with Petr
  • VMWare license vSphere follow up
  • IT email format follow up

next:
  • IT support while Stephen is away
  • 2019 plan, draft + input from team
  • phone interviews

PK

  • M3 demonstration
  • OVMS (waiting for the correct SD card to arrive)
  • TTL feedback and changes (configuring times, adding sounds)
  • ARGOS (TCS part seems to be done, as it was prepared and needs just to uncoment the code, working on MCSPU part)
  • Telemetry data for Xianyu
Next:
  • Collimation
  • ARGOS
  • OVMS after new card hopefully installed (Thursday?)

SH

  • mountain firewall upgrade completed
  • OVMS stuff with Petr
  • Whinging about R3 to all who listen.
  • IT #7345 -- right guy for Steward Saftety site is Dalton Porter (with MMT).
  • Clean up e-mail, IT for Doug Summers.
  • Steward e-mail help.

XZ

Last week:
  • Follow DX AdSec lab work;
  • OVMS+TT dicussion;
  • Support DX AdSec lab work, i.e, running script to start;
  • Get nessasry TCS telemetry from Petr and write an idl routine to extract data for AdSec gap analyse;
  • check the AOS code in git;
  • Read the MG electrinics design report.

This week:
  • finish read the MG electrinics design report;
  • support some other summer shutdown activities.

YZ

  • Collimation work:
    • Finished IT #7325, released an enhanced collimation data mining and searching tool. It allows us to extract any past collimation data with specific telescope instruments on either sides, bino/mono mode, and any time ranges.
    • Working on the next program on collimation, a replacement Python program for the Excel spreadsheet that will perform curve fitting of collimation data.
    • Work with Petr when he's back to study collimation data and procedures.
  • Collimation and pointing model documents.
  • M3 Retrofit training.
  • Continue to study the AO supervisor code and the IT issue #7261 (memory problem with WfsArbitrator).
Topic revision: r12 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback